OpenStudy (anonymous):

change 31/7 into a mixed #

OpenStudy (anonymous):

without using a calculator divide 31 by 7 and what is the remainder?

OpenStudy (anonymous):

its 4 right

OpenStudy (anonymous):

and the remainder?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

7*4 = 28 31-28 = 3 the remainder is 3

OpenStudy (anonymous):

so it would be 4 with a remainder of 3

OpenStudy (anonymous):

oh derr

OpenStudy (anonymous):

now, the 4 part is the whole number part of the mixed number the fraction part of the mixed number would be 3/7

OpenStudy (anonymous):

do you understand where the fraction part came from?

OpenStudy (anonymous):

yea i just got lost

OpenStudy (anonymous):

ok a mixed number consists of a whole number part, and a fraction yes?

OpenStudy (anonymous):

an example \[1 \frac{3}{4}\] 1 is the whole number and 3/4 is the fraction part

OpenStudy (anonymous):

\[\frac{31}{7} = \frac{28+3}{7} = \frac{28}{7} + \frac{3}{7} = 4 + \frac{3}{7}= 4 \frac{3}{7}\]